A confidence interval estimate for the population mean is given to be (39.50,47.85). If the standard deviation if 13.639 and the sample size is 41, answer each of the following (show all work):
A) Determine the maximum error of the estimate, E.
B) Determine the confidence level used for the given confidence interval.

Answer by stanbon(57413)

A confidence interval estimate for the population mean is given to be (39.50,47.85). If the standard deviation if 13.639 and the sample size is 41, answer each of the following (show all work):
A) Determine the maximum error of the estimate, E.
B) Determine the confidence level used for the given confidence interval.
---------------------------
The range of the confidence interval is always 2E.
----
2E = 47.85-39.50 = 8.35
E = 4.175
----
E = zs/sqrt(n)
4.175 = z*13.639/sqrt(41)
z = 1.96
So the confidence level is 95%
=================================
Cheers,
Stan H.
-----